OpenStudy (anonymous):

Divide using synthetic division P(x)= 9x5-27x3=81x-7 /3x+2

OpenStudy (anonymous):

is the prob like this? P(x)= 9x^5-27x^3=81x^-7 /(3x+2)

OpenStudy (anonymous):

yes it is .

OpenStudy (anonymous):

try cross multiplying it like this lets see what happen (9x^5-27x^3)(3x+2)=81x^-7

OpenStudy (anonymous):

yes we can lets try that

OpenStudy (anonymous):

okey, but i have no idea on how to solve fractions using synthetic division . I am slow learner when it comes in that situation . can u please help me ?

OpenStudy (anonymous):

ok here is the prob we have (9x^5-27x^3)(3x+2)=81x^-7 (9x^5-27x^3)(3x+2)81x^-7 =0 (9x^5-27x^3)(3x+2)81/x^7 =0

OpenStudy (anonymous):

is it also a way of getting the answers right ?but in the problem we have to use synthetic division. here is the problem can u help me to divide the fraction in synthetic as what i learn in synthetic u use multiplication, addition and subtraction. ok i will give you the problem then can u help me to solve it ? here it is :  9 0 0 -27 0 81 -7 how can we solve it in that manner ?

OpenStudy (anonymous):

hmm hold on we are not done yet now lets use synthetic division on it

OpenStudy (anonymous):

oh i've got mistake . here is the correct prob . sorry

OpenStudy (anonymous):

 here it is ..